OpenStudy (anonymous):

a girl delvering newspaper covers her route by traveling 2 locks west, 2 blocks north, then 9 blocks east. what is her direction? use counterclockwise as rhe positive angular direction from due east, between the limits of -180 egrees and 180 degrees. Her displacement is 7.28

OpenStudy (turingtest):

 add the vectors:the y-component is 2 the x-component is 2-9=-7 2/(-7)=tan (theta) theta=arctan (-2/7)=-0.278 radians we know from the picture that this is in the wrong quadrant (IV), so to get it into the proper quadrant (II) we add pi -0.278+(pi)=2.863 rad